OpenStudy (pgxgirl103):

check answer? what is the value of x? a. 68** b.62 c.112 d.124

OpenStudy (pgxgirl103):

I will attach a drawing- its not to scale

OpenStudy (redheadangel):

That right!

OpenStudy (anonymous):

This is an isosceles triangle.

OpenStudy (redheadangel):

correct**

OpenStudy (anonymous):

Sooo...the other angle in the bottom right will also be 56 degrees.

OpenStudy (anonymous):

o-o

OpenStudy (anonymous):

I'm just checking your answer aloud here....dunno why.

OpenStudy (anonymous):

56 + 56 = 112. 180 - 112 = 68...Yep, your answer is correct!

OpenStudy (pgxgirl103):

Thank you!

OpenStudy (igreen):

Yep, that's correct. The two angles at the bottom are congruent. 56 + 56 = 112. All 3 angles of all triangles add up to 180 so: 180 - 112 = 68